### v1

# Deep flows transmitted by forced surface gravity waves

Abstract
We examine a deep-water surface gravity wave packet generated by a pressure disturbance in the Lagrangian reference frame. The pressure disturbance has the form of a narrow-banded weakly nonlinear deep-water wave packet. During the forcing, the vorticity equation implies that the momentum is entirely in the near-surface Lagrangian mean flow. After the forcing is turned off, the wave packet propagates away from the forcing region carrying with it the energy imparted from the pressure forcing. These waves together with their induced long wave response have no momentum, in agreement with the classical results of Longuet-Higgins & Stewart (1962) and McIntyre (1981). The momentum imparted to the fluid from the form stress of the pressure forcing is entirely in a dipolar structure that remains in the area of the forcing event. This is in contrast with the finite-depth scenario discussed in McIntyre (1981) where the momentum imparted to the fluid by the forcing event propagates away as shallow water waves. Finally, we examine waves propagating from deep to shallow water and argue that wave packets, which initially have no momentum, may have non-zero momentum in finite-depth water through the trapping of a long wave at the step change in depth, which ensures the total momentum of the system remains conserved.
1. Introduction
Much of the ocean’s momentum is transferred from the atmosphere by form stresses pressing on the faces of surface gravity waves (Grare et al., 2013; Melville, 1996). Yet the fluid dynamical details of how atmospheric pressure is transmitted across the surface — somehow simultaneously forcing a current beneath the surface while also exciting a surface gravity wave — are mysterious. In this paper we elucidate some of these details through a detailed asymptotic analysis of the forced growth of a surface gravity wave packet of finite extent by an idealized atmospheric pressure distribution.
The unforced cousin of this problem was analyzed by Longuet-Higgins & Stewart (1962), who computed the total second-order mean flow associated with a freely-propagating deep-water surface gravity wave packet. Longuet-Higgins & Stewart (1962) showed that the surface-wave-induced mean flow has two components: first, a shallow surface component with no Eulerian-mean signature wholly attributable to the surface gravity wave Stokes drift, which may be evaluated diagnostically from and has similar vertical scales as the primary wave field. The second component of the mean flow is driven by the convergence and divergence of the primary flow, is purely Eulerian-mean, and is longer and deeper with scales similar to that of the wave packet. Remarkably and as emphasized by McIntyre (1981), the total vertically-integrated momentum of the freely-propagating packet vanishes at every horizontal point. In other words, the momentum carried by the deep return flow exactly cancels that of the shallow Stokes drift — implying that deep-water surface gravity waves do not carry momentum with them as they propagate.
The absence of momentum in freely-propagating wave packets begs the question: where is the momentum transferred from the atmosphere during wave generation? McIntyre (1981) argues that the momentum transferred during wave generation is carried away by packet-scale, shallow surface gravity waves — which being shallow, can carry momentum. But McIntyre (1981)’s answer, in addition to lacking the details that might accompany a more systematic solution of the equations of motion, produces counter-intuitive outcomes if the depth of the fluid is increased while holding the scale of the packet and atmospheric forcing fixed. For example, we find that atmospheric momentum would have to be instantaneously transferred farther and farther from the disturbance, carried by shallow water waves with longer and longer wavelengths without bound as the depth of the fluid increases.
This paper revisits this question through a detailed and systematic asymptotic solution of the Lagrangian equations of motion. Our aim is to elucidate the fate of atmospheric momentum transferred to a fluid during surface gravity wave packet generation, and in process, clarify McIntyre (1981)’s argument. We find, contrary to McIntyre (1981), that in an infinitely deep, non-rotating, and unstratified fluid, atmospheric momentum is not transferred to propagating waves but instead remains in place after forcing ceases while the packet propagates away. Moreover, we find the precise mechanism by which this occurs: as the packet grows, the pressure forces that drive its deep return flow are opposed — pointwise within the interior of the fluid — by the packet-scale component of the atmospheric pressure gradient. As a result, there is no deep return flow while the packet is forced, and it is only at the moment the forcing ceases that the deep return flow and its counterpart — the part that contains the net momentum transferred from the atmosphere — remains behind. We call this part of the flow, which is effectively transmitted from the surface into the deep ocean during the generation of surface gravity waves, and is given precisely as the negative of the deep return flow associated with the propagating packet at the moment the forcing ceases, as the “wave-transmitted flow”.
Our results are obtained through a multi-scale (in time and depth) asymptotic expansion of the equations of fluid motion formulated in a Lagrangian reference frame, which differs from usual analyses that leverage the Eulerian reference frame (Longuet-Higgins & Stewart, 1962; van den Bremer & Taylor, 2016; Haney & Young, 2017). In our Lagrangian analysis, the vorticity constraint — in our case, the vorticity of the fluid must vanish — plays a prominent role in identifying the shallow and deep components of the mean response to the atmospheric forcing (Salmon, 2020; Pizzo et al., 2023). This analysis highlights the close connection between the mean Lagrangian motion, the momentum of the fluid, and the vorticity of the fluid. As these flows are irrotational it might seem surprising that a vorticity analysis clarifies the dynamics – but indeed this perspective has proven to be useful in several contexts for water waves (Pizzo et al., 2023; Blaser et al., 2024) because in two dimensions vorticity is conserved on fluid particles, and so it is closely related to the mass flux.
An additional point of confusion arises from a close reading of Longuet-Higgins & Stewart (1962) and McIntyre (1981). Our analysis shows that when both the primary wave packet and the long wave response are in deep-water, neither carries momentum. Yet, shallow water waves can carry momentum, which seemingly leads to a violation of conservation of momentum if the wave packet propagates from deep to shallow water. To better understand this, we consider a wave packet in infinitely deep-water which experiences a step to finite depth and argue that this leads to the long wave being partially trapped at the step. When the momentum in each component is accounted for, the total momentum of the system remains conserved.
Figure 1. Contours of the streamlines in the Lagrangian reference frame for the second order mean flow of a deep-water surface gravity wave packet travelling from left to right, as indicated by the arrow above the wave packet. The colorbar indicates the value of the streamline contours, with blue being larger value than green. The velocity is sketched by the black arrows. The broad long wave response to the wave packet decays much more slowly with depth than the response to the carrier waves and is such that the momentum in the wave packet is exactly zero.
The outline of this paper is as follows. In §2 we present the equations of motion and several conservation laws for deep-water weakly nonlinear narrow banded surface gravity waves in the Lagrangian reference frame. In §3, we satisfy the equations of motion to third order in a small parameter. §4 considers the evolution of the waves and the mean flow after the forcing is turned off. The conclusions are presented in §6.
2. Equations of motion
We consider two-dimensional inviscid deep-water surface gravity waves. The Euler equations in Lagrangian coordinates are (Lamb, 1932, §13)
(1) $\mathcal{J}x_{\tau\tau}+p_{a}y_{b}-p_{b}y_{a}=0,$
(2) $\mathcal{J}y_{\tau\tau}+p_{b}x_{a}-p_{a}x_{b}+\mathcal{J}g=0,$
where $(x,y)$ are the particle locations, $(a,b)$ are particle labels, $p$ is the fluid pressure, $g$ is the acceleration due to gravity, and $\mathcal{J}$ is defined as
(3) $\mathcal{J}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}x_{a}y_{b}-x_{b}y_{a}.$
Derivatives with respect to the time $\tau$ take $(a,b)$ to be fixed, so that $\partial_{\tau}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}D/Dt$. Even though the acceleration term is linear in the Lagrangian frame, the pressure term is now nonlinear. This highlights the fact that the mapping between the two frames may be (highly) nonlinear.
The continuity equation in Lagrangian coordinates is
(4) 00 0 $=u_{x}+v_{y}\,,$
(5) $=\frac{\partial(x_{\tau},y)}{\partial(x,y)}-\frac{\partial(y_{\tau},x)}{\partial(x,y)}\,,$
(6) $=\left(\frac{\partial(x_{\tau},y)}{\partial(a,b)}+\frac{\partial(x,y_{\tau})}{\partial(a,b)}\right)\frac{\partial(a,b)}{\partial(x,y)}\,,$
which implies
(7) $\frac{\partial\mathcal{J}}{\partial\tau}=0\,.$
In order for the mapping to be invertible we require that $\mathcal{J}$ does not change sign. However, there are no other restrictions on $\mathcal{J}$, and this gauge invariance to how one labels the particles offers tremendously algebraic simplifications if the correct gauge is chosen. This symmetry also corresponds to the conservation of vorticity, $q$ (Salmon, 1988).
The vorticity $q$ is given by
(8) $q$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}v_{x}-u_{y}\,,$
(9) $=\frac{\partial(v,y)}{\partial(x,y)}+\frac{\partial(u,y)}{\partial(x,y)}\,,$
so that
(10) $q\mathcal{J}=\frac{\partial(x_{\tau},x)}{\partial(a,b)}+\frac{\partial(y_{\tau},y)}{\partial(a,b)},$
where the Euler equations in two-dimensions imply that
(11) $\frac{\partial q}{\partial\tau}=0.$
For irrotational flows, $q=0$.
In the absence of forcing, we require that the pressure $p$ vanish at the free surface, i.e. $p=0$ at $b=0$. Finally, we require that $y_{\tau}\to 0$ as $b\to-\infty$ so that there is no flow through the infinitely deep bottom.
The kinematic boundary condition is automatically satisfied in the Lagrangian reference frame. Although not pursued in detail here, the series convergence rates are faster in the Lagrangian frame (Longuet-Higgins & Cokelet, 1976; Clamond, 2007) than the Eulerian frame as particles tend to cluster in the crests.
2.1. Conservation laws
If we vertically integrate equation (1), we have
(12) $\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\mathcal{J}x_{\tau}{\rm d}b+\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}\right){\rm d}b=0.$
We can rewrite the terms involving the pressure by noting
(13) $\frac{\partial}{\partial a}\frac{1}{2}\int_{-\infty}^{0}\left(py_{b}-p_{b}y\right){\rm d}b=\frac{1}{2}\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}+py_{ab}-p_{ab}y\right){\rm d}b.$
Integration by parts implies
(14) $\frac{1}{2}\frac{\partial}{\partial a}\int_{-\infty}^{0}\left(py_{b}-p_{b}y\right){\rm d}b=\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}\right){\rm d}b+\frac{1}{2}\left(py_{a}-yp_{a})\right|_{b=0},$
so that the conservation of horizontal momentum becomes
(15) $\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\mathcal{J}x_{\tau}\ {\rm d}b+\frac{\partial}{\partial a}\int_{-\infty}^{0}py_{b}{\rm d}b=\frac{1}{2}\left(yp_{a}-py_{a})\right|_{b=0}.$
In the absence of pressure forcing at the surface, $p=p_{a}=0$ at the free surface so that the right hand side becomes zero. The second term here represents the momentum flux in this framework, which we define as $S$, where
(16) $S\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int_{-\infty}^{0}py_{b}\ {\rm d}b,$
so that
(17) $\frac{\partial\mathcal{I}}{\partial\tau}+\frac{\partial S}{\partial a}=\frac{1}{2}\left(yp_{a}-py_{a})\right|_{b=0},$
with $\mathcal{I}$ defined as
(18) =$\mathcal{I}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int_{-\infty}^{0}\mathcal{J}x_{\tau}{\rm d}b.$
In the absence of pressure forcing, we have
(19) $\frac{\partial\mathcal{I}}{\partial\tau}+\frac{\partial S}{\partial a}=0,$
so that the momentum is conserved.
We derive an energy integral by multiplying equation (1) by $x_{\tau}$ and equation (2) by $y_{\tau}$ and vertically integrating to find
(20) $\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\mathcal{J}\left(\frac{x_{\tau}^{2}+y_{\tau}^{2}}{2}+y\right){\rm d}b+\int_{-\infty}^{0}\left(p_{a}(x_{\tau}y_{b}-y_{\tau}x_{b})+p_{b}(y_{\tau}x_{a}-x_{\tau}y_{a})\right){\rm d}b=0.$
Now, we begin by considering the integral
(21) $\frac{\partial}{\partial a}\int_{-\infty}^{0}\left(p_{b}(xy_{\tau}-x_{\tau}y)+p(xy_{\tau b}-x_{\tau b}y)\right){\rm d}b.$
We may then show, by application of integration by parts that
(22) $\begin{split}&\int_{-\infty}^{0}\left(p_{a}(x_{\tau}y_{b}-y_{\tau}x_{b})+p_{b}(y_{\tau}x_{a}-x_{\tau}y_{a})\right){\rm d}b\\ &\qquad=\frac{\partial}{\partial a}\int_{-\infty}^{0}\left(p_{b}(xy_{\tau}-x_{\tau}y)+p(xy_{\tau b}-x_{\tau b}y)\right){\rm d}b+\left.\left(p_{a}(xy_{\tau}-x_{\tau}y)\right)\right|_{b=0},\end{split}$
so that
(23) $\frac{\partial E}{\partial\tau}+\frac{\partial F}{\partial a}=\left.\left(p_{a}(x_{\tau}y-xy_{\tau})\right)\right|_{b=0},$
where the energy is
(24) $E\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int_{-\infty}^{0}\mathcal{J}\left(\frac{x_{\tau}^{2}+y_{\tau}^{2}}{2}+y\right){\rm d}b,$
and the energy flux $F$ is given by
(25) $F\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int_{-\infty}^{0}p(x_{b}y_{\tau}-x_{\tau}y_{b}){\rm d}b.$
In the absence of pressure forcing the energy is conserved.
Next, consider the kinematic relationship $z_{\tau}^{\star}=\chi_{z}$, for $\star$ denoting complex conjugate and $\chi=\phi+\mathrm{i}\psi$ the complex velocity potential. If we then map to the Lagrangian frame, and let $s=a+\mathrm{i}b$, we have, $z_{\tau}^{\star}z_{w}=\chi_{w}$ (Longuet-Higgins, 1980b) so that the stream function $\psi$ in the Lagrangian reference frame satisfies
(26) $\psi_{a}=x_{\tau}y_{a}-y_{\tau}x_{a};\quad\psi_{b}=x_{\tau}y_{b}-y_{\tau}x_{b}.$
From this association, we can then write the energy flux as
(27) $F=\int_{-\infty}^{0}p\psi_{b}{\rm d}b.$
2.2. Momentum and energy transfer by surface pressure forcing
We now consider the total momentum conservation by integrating (15) over $a$ and assuming that the pressure is compact so that upon integrating by parts we find
(28) $\frac{\partial}{\partial\tau}\iint\mathcal{J}u{\,\rm d}a{\,\rm d}b=\int py_{a}\,|_{b=0}{\,\rm d}a\,.$
The left side of (28) is the rate of change of total horizontal momentum. The right side of (28) is the form stress: the only way by which momentum enters an inviscid fluid in a closed system.
Similarly, we can look at the dynamical budget. Once again, we average over all horizontal particle labels to find
(29) $\frac{\partial}{\partial\tau}\iint\mathcal{J}\left(\frac{x_{\tau}^{2}+y_{\tau}^{2}}{2}+y\right){\,\rm d}a{\,\rm d}b=\int\left.p(x_{\tau}y_{a}-x_{a}y_{\tau})\right|_{b=0}{\rm d}a.$
The right hand side of the momentum balance is a term related to the form drag, while the right hand side of the energy equation is related to the spin and angular momentum of the fluid (Longuet-Higgins, 1980a).
3. Weakly nonlinear narrow-banded waves
Our goal is to find $(x,y,p)$ to third order in the small parameter $\epsilon\ll 1$. First, we satisfy conservation of mass and vorticity. We then find an expression for the pressure $p$. Finally, the dispersion relationship is found by ensuring that the pressure condition is met at $b=0$.
To solve our system we expand $x,y,p$ in asymptotic series,
(30) $x=\sum_{n}\epsilon^{n}x^{(n)};\quad y=\sum_{n}\epsilon^{n}y^{(n)};\quad p=\sum_{n}\epsilon^{n}p^{(n)},$
where $x^{(n)}=x^{(n)}(a,b,\tau,\epsilon a,\epsilon b,\epsilon\tau,...),y^{(n)}=y^{(n)}(a,b,\tau,\epsilon a,\epsilon b,\epsilon\tau,...),p^{(n)}=p^{(n)}(a,b,\tau,\epsilon a,\epsilon b,\epsilon\tau,...)$. Here, the superscript with parentheses represents the $n-th$ term in the expansion, not a power of that variable.
In our analysis we take $\mathcal{J}=1$, which amounts to a gauge choice (Salmon, 2020) sometimes known as Miche’s condition (Clamond, 2007).
The pressure forcing $p^{\mathrm{atm}}$ is defined at the free surface, $b=0$, at first order to be
(31) $p^{\mathrm{atm}}=2g\epsilon\mathcal{A}\sin\theta,$
where the phase $\theta$ is defined
[…]
and $\mathcal{A}=\mathcal{A}(\epsilon a,\epsilon\tau)$ is a slowly varying function (with units of length), which represents the modulation to the forcing event in space and time.
3.1. Zeroth order expansion
At zeroth order, the system is in hydrostatic balance and we have
(33) $x^{(0)}=a;\quad y^{(0)}=b;\quad p^{(0)}=-gb.$
3.2. First order expansion
To first order in $\epsilon$, conservation of mass requires
(34) $x^{(1)}_{a}+y^{(1)}_{b}=0,$
while the vorticity equation implies
(35) $-x^{(1)}_{\tau b}+y^{(1)}_{\tau a}=0.$
That is, the first order fluid velocity $(x^{(1)}_{\tau},y^{(1)}_{\tau})$ satisfies the Cauchy-Riemann equations, with each component being harmonic with respect to $(a,b)$.
The $O(\epsilon)$ momentum equations are
(36) $u^{(1)}_{\tau}+p^{(1)}_{a}+gx^{(1)}_{b}=0\,,$
(37) $v^{(1)}_{\tau}+p^{(1)}_{b}+gy^{(1)}_{b}=0\,.$
[…]
(38) $\triangle p^{(1)}=0\,.$
(38) has the surface boundary condition $p^{(1)}\,|_{b=0}=p^{\mathrm{atm}}=2g\mathcal{A}\sin\theta$. The fluid has infinite depth so that we apply the condition that $p^{(1)}\,|_{b\to-\infty}\to 0$. The solution that satisfies (38), $p^{(1)}=2g\mathcal{A}\sin\theta$ at $b=0$ and $p^{(1)}\to 0$ as $b\to-\infty$ is
(39) $p^{(1)}=2g\mathcal{A}\sin\theta e^{kb}$
where we have assumed that $k>0$, that is all waves are travelling to the right in our system.
We differentiate the momentum equations once more with respect to time, and use the Cauchy-Riemann equations to find
(40) $u^{(1)}_{\tau\tau}+gu^{(1)}_{b}=-p^{(1)}_{\tau a},$
(41) $v^{(1)}_{\tau\tau}+gv^{(1)}_{b}=-p^{(1)}_{\tau b}.$
Taking $(u^{(1)},v^{(1)})$ to be separable with vertical dependence $e^{kb}$ implies
(42) $u^{(1)}_{\tau\tau}+gku^{(1)}=-p^{(1)}_{\tau a},$
(43) $v^{(1)}_{\tau\tau}+gkv^{(1)}=-p^{(1)}_{\tau b}.$
This is the governing equation for a (resonantly) forced harmonic oscillator. Taking the initial conditions to be that the surface is initially flat, i.e. $(x^{(1)},y^{(1)})=(0,0)$ at $\tau=0$, this has solution
(44) $x^{(1)}=-\frac{\mathcal{A}\omega\tau}{2}\cos\theta e^{kb}$
(45) $y^{(1)}=\frac{\mathcal{A}\omega\tau}{2}\sin\theta e^{kb}.$
The free surface is growing secularly in time to first order in $\epsilon$. By forcing the system with a wave form that exactly solves the equations of motion at this order, we eliminate the need to resolve dispersive corrections (c.f. the Cauchy-Poisson problem) that would unnecessarily complicate the analysis.
3.3. Second order expansion
To second order in $\epsilon$, we find that mass conservation implies
(46) $x^{(2)}_{a}+y^{(2)}_{b}=\omega^{2}\tau^{2}k^{2}\mathcal{A}^{2}e^{2kb}$
and vorticity conservation implies
(47) $x^{(2)}_{\tau b}-y^{(2)}_{\tau a}=2\omega\tau^{2}k^{2}\mathcal{A}^{2}e^{2kb}.$
By differentiating (46) with respect to time, these equations can be reduced to
(48) $\Delta u^{(2)}=4\omega^{3}\tau^{2}k^{3}\mathcal{A}^{2}e^{2kb},$
(49) $\Delta v^{(2)}=4\omega^{2}\tau k^{3}\mathcal{A}^{2}e^{2kb}.$
Unlike in the Eulerian frame, at second order these equations have no wavy terms and therefore represent the mean flow response to the pressure forcing.
We are interested in solving for the behavior of $u^{(2)}$, so we must find a boundary condition on this term. The horizontal momentum equation implies that at second order
(50) $u^{(2)}_{\tau}=-p^{(2)}_{a}-gy^{(2)}_{a}+2g\omega\tau k^{2}\mathcal{A}^{2}.$
Taking a spatial average, $(2\pi)^{-1}\int_{0}^{2\pi}{\rm d}a$, at $b=0$ this reduces to
(51) $u^{(2)}=\omega\tau^{2}k\mathcal{A}^{2},$
while we also require $u^{(2)}\to 0$ as $b\to-\infty$.
Equation (48) with boundary condition (51) has solution
(52) $u^{(2)}=\frac{\omega}{k}k^{2}\mathcal{A}^{2}\omega^{2}\tau^{2}e^{2kb}.$
The mean momentum of the system is
(53) $\left\langle\int_{-\infty}^{0}\mathcal{J}x_{\tau}{\rm d}b\right\rangle=\epsilon^{2}\int_{-\infty}^{0}\mathcal{J}u^{(2)}{\rm d}b=\epsilon^{2}\frac{\omega}{2k}k^{2}\mathcal{A}^{2}\omega^{2}\tau^{2}+O(\epsilon^{4}),$
where the angle brackets are a spatial average over one period. This integral is non-zero and implies that the momentum is entirely contained in the drift of the carrier waves. Note, the phase averaged momentum is conserved in a local sense at each horizontal location (in the slow variable $\epsilon a$), as well as in a global sense when integrating over all of $a$. To interpret these results it is instructive to juxtapose this with the unforced scenario.
Figure 2. The streamlines for the second order mean flow of a forced deep-water surface gravity wave packet travelling from left to right, as indicated by the arrow above the wave packet. During forcing there is no long wave response to the wave packet. This is interpreted as the forcing generating an equal and opposite return flow that exactly cancels the long wave, resulting in their only being the carrier wave drift in the forcing region. This is shown below the dashed line.
4. Behavior after forcing is turned off
We now turn off the forcing at a time $\tau=T$, allowing us to follow Salmon (2020) closely. At first order,
(54) $x^{(1)}=-\omega T\mathcal{A}e^{kb}\cos\theta,$
(55) $y^{(1)}=\omega T\mathcal{A}e^{kb}\sin\theta,$
(56) $p^{(1)}=0,$
and $\omega^{2}=gk$. These waves are no longer growing in time and represent linear deep-water surface gravity waves. At second order, the equations governing the mean flow become
(57) $x^{(2)}_{a}+y^{(2)}_{b}=\omega^{2}T^{2}k^{2}\mathcal{A}^{2}e^{2kb}$
and vorticity conservation implies
(58) $x^{(2)}_{\tau b}-y^{(2)}_{\tau a}=2\omega T^{2}k^{2}\mathcal{A}^{2}e^{2kb}.$
Differentiating the mass conservation equation with respect to time, we have
(59) $x^{(2)}_{\tau a}+y^{(2)}_{\tau b}=0+O(\epsilon^{3}),$
which implies we can define a stream function for the second order mean flow:
(60) $x^{(2)}_{\tau}=-\psi_{b};\quad y^{(2)}_{\tau}=\psi_{a}.$
The equation governing the mean flow becomes
(61) $\Delta\psi=-2k^{2}\omega^{3}T^{2}\mathcal{A}^{2}e^{2kb}.$
Next, the pseudomomentum is defined as
(62) $\mathbb{P}=\mathbb{k}\tilde{\mathcal{A}}=\mathbb{k}E/\omega,$
where $\tilde{\mathcal{A}}=\omega T\mathcal{A}$. Therefore, the right hand side of (61) is the curl of the pseudomomentum:
(63) $\boldsymbol{\nabla_{a}}\times\mathbb{P}=-2k^{2}\omega^{3}T^{2}\mathcal{A}^{2}e^{2kb}\hat{z},$
where $\boldsymbol{\nabla_{a}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(\partial_{a},\partial_{b})$ and $\hat{z}$ points out of the page.
The boundary condition for $\psi$ is found once again from the horizontal momentum equation and is $\psi=0$ at $b=0$. This immediately implies that the total momentum is zero, as
(64) $\frac{1}{2\pi}\int_{0}^{2\pi}\int_{-\infty}^{0}\mathcal{J}x_{\tau}{\rm d}b{\rm d}a=\psi(b=0)+O(\epsilon^{3}).$
We can consider the particular and homogeneous solution to this system. That is, let $\psi=\psi_{c}+\psi_{LW}$ where $c$ denotes carrier and $LW$ the long wave contribution to the stream function. By inspection,
(65) $\psi_{c}=-\omega^{3}T^{2}\mathcal{A}^{2}e^{2kb},$
agreeing with the classical prediction of Stokes. This implies $\psi_{LW}$ must obey the relationship
(66) $\Delta\psi_{LW}=0$
with boundary condition $\psi_{LW}(b=0)=-4\omega^{3}T^{2}\mathcal{A}^{2}$. The domain of the labels here is the lower half plane, so that we may use the theory of Poisson kernels to write down the solution
(67) $\psi_{LW}=-\frac{4\omega^{3}T^{2}k^{2}}{\pi}\int_{-\infty}^{\infty}\frac{\mathcal{A}^{2}b}{(a^{\prime}-a)^{2}+b^{2}}{\rm d}a^{\prime}.$
$\psi_{c}$ and $\psi_{LW}$ are used to generate the figures in this manuscript. Note, some care must be taken when considering these functions at $b=0$, as $\psi_{LW}$ is not zero there, but instead reduces to the prescribed boundary condition with $b(1+(a-a^{\prime})^{2}/b^{2})^{-1}$ being a $\delta-$function sequence as $b\to 0$.
The depth dependence of the long wave response is an integral of algebraic functions, as opposed to the exponential decay of the carrier waves. This is reflected in the fact that the contours in figure 1 near the surface have a much larger gradient with respect to label $b$ than the long wave response. In the limit that the amplitudes are uniform except at the edges of the packet, we return to the image presented in McIntyre (1981).
The energy is still entirely contained in wave motion – that is the return flow does not contribute to this budget at this order. Similarly, one can show that the momentum flux, $S$ as defined by the phase average of (16), is entirely supported by the carrier wave motion and has no contribution form the long wave response.
The above analysis corroborates the computations by Longuet-Higgins & Stewart (1964) and McIntyre (1981) and offers a path towards interpreting the results from the forcing case. As the mean flow is entirely in the carrier wave drift during forcing, and the mean flow is linear at this order, we conclude that there is an equal and opposite deep flow that exactly cancels the return flow. This scenario is shown in figure 2. The forcing is subsequently turned off and the wave packet propagates away from the generation region. This is displayed in figure 3. The packet and its return flow have no momentum. In the forcing region, we are left with the dipole that was created during the forcing event. As this can only propagate under its own image (chosen to ensure that the free surface acts as a rigid lid for this flow), this will be extremely weak and the momentum is effectively localized to this forcing region.
Figure 3. Left: The second order mean stream lines during the forcing event. The momentum resides entirely in the carrier wave drift. During the forcing event, the long wave response is entirely cancelled by an equal and opposite flow induced by the forcing. Right: the forcing is now turned off and the waves propagate away from the forcing region. The wave packet again has zero net momentum, agreeing with the classical result. In the region of forcing, a dipolar flow containing the momentum imparted by the forcing event.
What happens as these waves approach shore? As was shown in (64), unforced deep-water wave packets have no momentum. However, Longuet-Higgins & Stewart (1964) showed that if the carrier waves are still in deep-water, but the packet is in finite depth, the wave packet possesses momentum which we denote $I_{0}$. If a wave packets starts off so that the waves and the packet are in infinitely deep-water and then impinge upon a step such that the depth changes so that the wave packet is now in finite depth, this would seem to violate conservation of momentum as the packet has experienced a change in momentum, $I_{0}$. To balance this, somewhere in the fluid there must be momentum with value $-I_{0}$. As the carrier waves are still in deep-water, the drift (and hence momentum) associated with the carrier waves does not change over the step. Therefore, the long wave is adjusting, and weakening, when it is in finite depth versus deep-water. That excess momentum is jettisoned at the step. As it can only advect weakly under its own image charge, this flow is essentially bound to the step region.
5. Conclusion
We find that during the forcing event, the fluid has momentum which is coincident with the Stokes drift of the waves. Moreover, the additional momentum that would be associated with the deep return flow is precisely cancelled by deep interior pressure gradients sustained by the atmospheric pressure forcing at the surface. But we also find, like Longuet-Higgins & Stewart (1962), that the Stokes drift and its associated deep Eulerian return flow are carried away from the generation region with the wave packet as soon as the forcing ceases and the packet begins to freely propagate. This leads to the remarkable conclusion that ultimately, the momentum transferred from the atmosphere — which is transmitted by deep pressure forces during wave generation — resides in a deep mean flow that identically opposes the deep Eulerian return flow up to the moment that forcing stops. In non-rotating, unstratified deep water, this mean flow is non-propagating and thus remains in place as the new surface gravity wave packet propagates away.
Besides the modification to the momentum budget considered here, Dysthe (1979) argued that the inclusion of the long wave mean flow modifies the modulation instability growth rates and it was subsequently shown that this gives more accurate predictions that the lower order nonlinear Schrodinger equation (Melville, 1982). However, during forcing this long wave is not presented implying a modulation to the instability growth rates.
A natural question is whether stratification or rotation will significantly change the picture presented here. If the atmospheric forcing is nearly impulsive, we need consider only the initial value problem in which a surface gravity wave packet is superimposed with a Lagrangian-mean flow given by the Stokes drift, without a deep Eulerian return flow — the state of the system at the conclusion of an impulsive atmospheric forcing resonant with a surface gravity wave packet. In a stratified fluid, the resulting evolution of the freely-propagating packet was analyzed by Haney & Young (2017), who showed that the deep Eulerian return flow is slightly modified by the presence of realistic oceanic stratfication and in addition weakly radiates internal waves as the packet propagates. The mean flow left behind by the atmospheric forcing is thus accordingly weakly modified by stratification, in the opposite manner, and likely also decomposes into a weak propagating internal response together with a stronger stationary vortical response. Less clear is the fate of atmospheric momentum and evolution of a freely-propagating packet in a rotating fluid. Unlike stratification, rotation likely dramatically changes the induced flow surrounding a freely-propagating packet. For example, Ursell & Deacon (1950) (see also Pollard, 1970) showed that the shallow, Stokes-drift-associated part of the wave-induced mean flow cannot persist in a rotating fluid. As a result, the circulation around a freely-propagating packet in a rotating fluid is likely much weaker. We may then speculate that the flow left behind by the packet, which contains the momentum transferred from the atmosphere, takes the form of a packet-scale near-inertial wave with an initial state given by the wave packet Stokes drift profile.
We thank Aidan Blaser and T.S. van den Bremer for useful comments. N.P. was partially supported by NSF OCE, NASA S-MODE and NASA OVWST.